OpenStudy (anonymous):

. Charlie measured the circle-shaped part of a sun his sister drew on the chalkboard. The circumference of the circle is 75.36 cm. What is the diameter? (Use 3.14 for .)

Nnesha (nnesha):

do you know the formula ??

OpenStudy (anonymous):

no

Nnesha (nnesha):

\[\huge\rm C = \pi \times d\] C =circumference d = diameter replace variables by their values and solve or missing variable

Nnesha (nnesha):

for*

OpenStudy (anonymous):

i dont get it

Nnesha (nnesha):

circumference of the circle is what ??

OpenStudy (anonymous):

3.14?

Nnesha (nnesha):

nope read the question again

OpenStudy (anonymous):

o 75.36 sorry

Nnesha (nnesha):

C= 75.36 pi = ?? d = ?? you have to solve for d so replace pi and c by their values

OpenStudy (anonymous):

pi is 3.14?

Nnesha (nnesha):

yep right

OpenStudy (anonymous):

so it is pi times the ....um diameter

Nnesha (nnesha):

c = pi times diameter how would you substitute ??

OpenStudy (anonymous):

i dont kno a variable?

Nnesha (nnesha):

okay that's right but i'm saying how would equation look like when you replace c and pi by their values

OpenStudy (anonymous):

o 75.36 = 3.14 * d

Nnesha (nnesha):

yep right so how would you solve for d? do you know ??

OpenStudy (anonymous):

no i dont

Nnesha (nnesha):

3.14 is multiplying by d do cancel out 3.14 from right side you should do opposite of multiply which is ?

OpenStudy (anonymous):

divide

Nnesha (nnesha):

yep right so divide both side by 3.14

OpenStudy (anonymous):

24 and 1

Nnesha (nnesha):

yep right d times 1 = d

OpenStudy (anonymous):

so it is 24? rite

Nnesha (nnesha):

yes

OpenStudy (anonymous):

wait but this guy said 36 r6 2 = 6 how did he get 6?

OpenStudy (anonymous):

r^2*

Nnesha (nnesha):

36 times r^2 = 6 ??

OpenStudy (anonymous):

yes that is wvat he said

Nnesha (nnesha):

2 times 36 times r^2 ?? write one more time i don't get that

OpenStudy (anonymous):

he said that 36= r^2 and so radius =6

Nnesha (nnesha):

is that a same question or different ?

OpenStudy (anonymous):

different

Nnesha (nnesha):

okay so r = d/2 radius is half of diameter r^2 = 36 <--- how would you solve this for r ??

OpenStudy (anonymous):

um 36/2?

Nnesha (nnesha):

nope r^2 = 36 this one

Nnesha (nnesha):

you have to take square root both side

OpenStudy (anonymous):

square root?

Nnesha (nnesha):

yep  you can change square root to exponent \[\huge\rm \sqrt[n]{x^m} = x^\frac{ m }{ n }\]